OpenStudy (anonymous):

\[\int\limits_{}^{}\int\limits_{D}^{}y ^{2}e ^{xy}dA\]D is bounded by y=x. y=4, x=0 Set up iterated integrals for both orders of integration. Then evaluate the integral using the easier order and explain why it is easier.

OpenStudy (anonymous):

I know I can do... \[\int\limits_{0}^{4}\int\limits_{x}^{4}y ^{2}e ^{xy}dydx\]but I'm not sure what they mean by both orders of integration and how to find the other orders.

OpenStudy (turingtest):



OpenStudy (turingtest):

\[x\le y\le4,~~0\le x\le4\implies0\le x\le y,~~0\le y\le4\]

OpenStudy (turingtest):

does that make any sense to you?

OpenStudy (anonymous):

Yep, I get that.

OpenStudy (turingtest):

so if you switch the order, those would be your bounds where are you stuck?

OpenStudy (anonymous):

How would I break up the integral into iterated sections since the x is in e^(xy)?

OpenStudy (turingtest):

\[\int_0^4\int_0^yy^2e^{yx}dxdy=\int_0^4y\left[\int_0^y ye^{yx}dx\right]dy\]

OpenStudy (turingtest):

notice that\[\frac{\partial}{\partial x}e^{yx}=ye^{yx}\]

OpenStudy (anonymous):

Alright, so I would have.. \[\int\limits_{0}^{4}ydy*[e ^{xy}|0 \ \to\ y]?\]

OpenStudy (turingtest):

yes

OpenStudy (anonymous):

Thank you!

OpenStudy (turingtest):

welcome :D